# Active Learning with Task–Driven Representations for Messy Pools

### v1

@@ 4.2 A Representation Fine–Tuning Approach
The approach naturally starts with some initial representation defined by an encoder $g:X\to\mathbb{R}^{d}$. This can either be trained directly on the pool data using any powerful unsupervised representation learning technique—such as those based on contrastive learning (e.g. SimCLR (Chen et al., 2020a, )), clustering (e.g. SwAV (Caron et al.,, 2020), DeepCluster (Caron et al.,, 2018)) or masked autoencoders (e.g. MAE (He et al.,, 2022))—or it can be taken as a fixed pretained encoder trained on some other data, such as ESM-3 for protein sequences (Hayes et al.,, 2025). In the latter case, the representation does not necessarily need to have been trained in an unsupervised manner itself, but it will inevitably not have information about the labels of the task at hand as these are yet to be collected. For our experiments we will use an encoder trained on the pool data using SimCLRv2 (Chen et al., 2020b, ).
Once initialised, we extend $g$ by adding a guidance classifier $c_{\omega}:\mathbb{R}^{d}\to[0,1]^{|\mathcal{Y}|}$ to its final layer, resulting in the model $g\circ c:X\to[0,1]^{|\mathcal{Y}|}$, which we train in a fully supervised manner using the acquired labels following (Chen et al., 2020b, ). The final representations are taken from $g$ after this fine–tuning as before. For $c_{\omega}$, we follow Chen et al., 2020b and use what is effectively a 1-layer neural network with a modest number of hidden units. For the same reasons discussed in Section 4.1, we do not use $c_{\omega}$ for our prediction head, again using a random forest in our experiments instead. Further algorithmic details on the precise approach used for the experiments and ablations with different prediction heads are given in the Appendix.
Table 1: Final test accuracy of different active learning methods on the F+MNIST, CIFAR-10+100 and CheXpert datasets. We report the mean $\pm 1$ standard error over 4 seeds. The full active learning curves given in the Appendix.
@@
SIMILAR (Kothawade et al.,, 2021) $93.82\pm 0.18$ $30.87\pm 1.57$ $71.94\pm 0.47$
GALAXY (Zhang et al.,, 2022) $84.74\pm 0.74$ $55.28\pm 0.42$ $78.76\pm 0.35$
Cluster Margin (Citovsky et al.,, 2021) $94.24\pm 0.17$ $32.79\pm 0.45$ $75.41\pm 0.29$
US+EPIG (SimCLRv2, Bickford Smith et al., (2024)) $98.53\pm 0.12$ $6.19\pm 0.42 77$$7.84\pm 0.28$
US+EPIG (VAE, Bickford Smith et al., (2024)) $94.50\pm 0.34$ $30.47\pm 1.17$ $66.69\pm 0.56$
@@ 5 Related Work
Previous works on active learning with messy, uncurated pools have primarily focused on developing new acquisition strategies (Zhang et al.,, 2022; Kothawade et al.,, 2021; Zhang et al.,, 2023; Nuggehalli et al.,, 2023; Russakovsky et al.,, 2015), neglecting the importance of using a model that incorporates information from the unlabelled data in a way that accounts for the messiness of the pool. Indeed, they have mainly used fully supervised models, with some approaches initialising their model with weights pre–trained on ImageNet in a fully supervised fashion or from foundations models such as CLIP (Radford et al.,, 2021; Xie et al.,, 2024; Nuggehalli et al.,, 2023). We compare with these alternatives in Section 6.2 and find they significantly underperform compared to using representations trained on the pool.
@@
On the other hand, various work have considered using semi–supervised models in active learning (Burkhardt et al.,, 2018; Seo et al., 2022b, ; Osband et al., 2022a, ; Mittal et al.,, 2023; Sener and Savarese,, 2017; Yehuda et al.,, 2022; Lüth et al.,, 2023), with several works showing supervised models significantly lagging behind semi–supervised ones in terms of downstream performance (Bickford Smith et al.,, 2024; Hacohen et al.,, 2022; Yehuda et al.,, 2022). In general, there have been questions raised about the benefits of active learning when semi–supervised models are used: many acquisition strategies designed for fully supervised models having been shown to no longer provide reliable gains over random acquisition in the semi-supervised setting (Sener and Savarese,, 2017; Yehuda et al.,, 2022; Lüth et al.,, 2023; Chan et al.,, 2021; Bhatnagar et al.,, 2020; Ebrahimi et al.,, 2020; Gao et al.,, 2020). However, Bickford Smith et al., (2024) recently showed that EPIG does reliably outperform random acquisition, as well as various other acquisition strategies, in this setting, using their unsupervised representation learning approach.
@@
6 Experiments
We refer to the approaches introduced in Sections 4.1, 4.2 as TD-SPLIT and TD-FT respectively. To validate them, we first compare with various baselines that have been specifically designed for, or shown strong performance on, the messy pool scenarios we are interested in (c.f. §3). As part of this, we include different variants of the unsupervised representation (US) approach of Bickford Smith et al., (2024). All experiments were run on an NVIDIA H100 80GB GPU.
Datasets:
To construct datasets with redundant classes and class imbalance, we combined existing benchmarks. Our F+MNIST, uses the digits “5” and “6” from MNIST (Deng,, 2012) as the target classes for active learning, while the entire FashionMNIST (Xiao et al.,, 2017) dataset is included as redundant data. CIFAR-10+100, uses the first five classes of CIFAR-10 (Krizhevsky,, 2009) as its target classes, with the full CIFAR-100 dataset (Krizhevsky,, 2009) serving as the redundant data.
We further use the CheXpert (Irvin et al.,, 2019) dataset as an example of a real–world dataset with redundant information and existing class imbalance. CheXpert comprises of chest X–rays take from a variety of patients from different angles. We consider the binary classification task of identifying pleural effusion, i.e. fluid in the corner of the lungs. The redundant features here are the many anatomical and acquisition–related variations (e.g., bones, implants, soft tissue, scan artefacts) that are irrelevant to the diagnosis and generally represent larger image variations than the target–relevant information (Joy et al.,, 2021).
Representation Learning: For all datasets, we use VAE (Kingma et al.,, 2013) and SimCLRv2 (Chen et al., 2020b, ) encoders, pairing them with TD–SPLIT and TD–FT respectively. We also include unsupervised variants following Bickford Smith et al., (2024) (noting they themselves also consider the same two encoder strategies). For all approaches, we adopt the learning rate, batch size, and optimizer from the original papers and train each model for 500 epochs.
Models: For F+MNIST and CheXpert, we use the Burgess encoder (Burgess et al.,, 2017) (and decoder) for our VAE–based representations and otherwise use a ResNet18 encoder (He et al.,, 2016). For CIFAR-10+100, we replace the ResNet18 with a ResNet50 and the Burgess encoder with a ResNet–VAE (Kingma et al.,, 2016).
Active learning: We run active learning for a budget of 500, 6000, and 10,000 labels for F+MNIST, CheXpert and CIFAR–10+100 respectively. We use batch acquisitions for all datasets, with a batch size of 10 for F+MNIST and 100 for the others. We use the “power” batch acquisition strategy from Kirsch et al., (2021) with $\beta=4$ for F+MNIST and $\beta=8$ otherwise. We make this choice as this strategy is both highly scalable and has been shown to give performance comparable to more sophisticated batch acquisition strategies. We re–train our semi–supervised encoders every 5 acquisition rounds and ablate with different re–training periods in the Appendix.
@@ 6.1 Comparison with Existing Approaches
To highlight the effectiveness of our approach, we first compare against baselines designed for (or shown good performance on) problems with messy pools: SIMILAR (Kothawade et al.,, 2021), Cluster Margin (Citovsky et al.,, 2021), and GALAXY (Zhang et al.,, 2022), and the approach in Bickford Smith et al., (2024) which we refer to as US+EPIG. We also compare with random acquisition using the US, TD-SPLIT, and TD-FT approaches. We implement the baselines as in the original papers.
@@
Our TD-SPLIT method also shows strong performance relative to the baselines, albeit generally underperforming TD-FT. This is because it is using a much more lightweight encoder: it still consistently and comprehensively outperforms the analogous US+EPIG approach with VAE–based representations, thus again emphasising the important of using task–driven representations.
@@
Figure 3: Test accuracy on F+MNIST using the US approach and our task–driven approach. Top row shows the results using VAE–based encoders and the bottom row shows the results for SimCLRv2 encoders. Experiments run for 4 seeds.
@@
Figure 4: Test accuracy on CheXpert using the US approach and our task–driven approach. Top row shows the results using VAE–based encoders and the bottom row shows the results for SimCLRv2 encoders. Experiments run for 4 seeds.
@@ 6.2 Pool-Based Task-Driven Representations Outperform Transfer Learning Based AL
A common setup in AL papers that deal with messy pools is a transfer–learning approach where a fully supervised model initialised from a large pretrained/foundation model is fine–tuned to the task (Zhang et al.,, 2022, 2023; Xie et al.,, 2024; Nuggehalli et al.,, 2023; Bommasani et al.,, 2021; Gupte et al.,, 2024). Our TD-FT framework can leverage these models by initialising the encoder, $g$, with their pre–trained weights instead of using an unsupervised representation of the pool. We test this on the CIFAR-10+100 dataset by comparing our standard TD-FT method against instead initializing the encoder using both publicly available pre-trained models—namely ResNet-50 and ResNet-101 pre–trained with supervision on ImageNet (Russakovsky et al.,, 2015)—and foundation models—namely two variants of the self-supervised DINOv2 model (Oquab et al.,, 2023), pre–trained on a large, curated dataset of 142 million unlabelled images.
@@ 6.3 Alternative Acquisition Strategies
As an ablation to demonstrate the broader benefits of our task–driven approach with other acquisition strategies, we apply our TD-FT and TD-SPLIT approaches to BALD (Houlsby et al.,, 2011) and Confidence Sampling (Settles,, 2009), comparing to analogous unsupervised representation (US) setups.
@@ Acknowledgements
Supplementary Materials for Active Learning with Task–Driven Representations for Messy Pools
Supplementary Contents
@@ A.1.1 Datasets
We used the F+MNIST as an example of a dataset with redundant classes and class imbalance by combining existing benchmarks. Specifically, we used the digits “5” and “6” from MNIST (Deng,, 2012) as the target classes for active learning while the entire FashionMNIST dataset (Xiao et al.,, 2017) was included as the redundant data. We used an imbalance ratio of 10 for our pool, where the minority classes were chosen to be our target classes.
@@
We used the CIFAR-10+100 as another example of a dataset with redundant classes and class imbalance by combining existing benchmarks. Specifically, we used the first 5 classes of CIFAR-10 (Krizhevsky,, 2009) as our target classes for active learning while the entire CIFAR-100 dataset (Krizhevsky,, 2009) was included as the redundant data. We again used an imbalance ratio of 10 in our pool, where the minority classes were chosen to be our target classes.
@@
We used the CheXpert (Irvin et al.,, 2019) dataset as an example of a real–world dataset with redundant information and existing class imbalance. CheXpert comprises of chest X–rays taken from a variety of patients from different angles. We considered the binary classification task of identifying pleural effusion, i.e. fluid in the corner of the lungs. We filtered out observations with “NA” as the response for the pleural effusion task. The imbalance ratio in our pool was 2.5.
@@ A.1.2 Representation learning
For all datasets, we used VAE (Kingma et al.,, 2013) and SimCLRv2 encoders, pairing them with TD-SPLIT, TD-FT respectively. We also used their unsupervised variants for the US approach (Bickford Smith et al.,, 2024), where we first encoded our data points into the latent space of the unsupervised encoder, then performed active learning on the latent space with a prediction head. Below, we describe the details of the representation learning methods used.
@@
For TD-SPLIT, we followed the CCVAE (Joy et al.,, 2021) approach described in Section 4.1 and optimised objective (1). We chose $\alpha$ so that the labeled loss roughly matched the scale of the unlabeled loss. We performed the optimisation using stochastic gradient ascent where updates with the labelled and unlabelled data were conducted in separate batches. As the labelled dataset contained far fewer data points, we evenly spaced the labelled batches between the unlabelled ones. We trained our model for 500 epochs and, following (Joy et al.,, 2021), we used a batch size of $200$ for both the unlabeled and labeled data, the Adam optimiser (Kingma and Ba,, 2014) and a learning rate of $2\times 10^{-4}$. Table 2 shows the sizes for $z$, $z_{c}$ and values of $\alpha$.
For the US approach, we followed Burgess et al., (2017) and optimised their ELBO objective with $\beta=1$ (this amounts to the standard VAE objective). We used the Adam optimiser, a learning rate of $5\times 10^{-4}$, a batch size of $200$ and KL annealing as in Burgess et al., (2017). For all datasets, we used the same latent dimensions as for TD-SPLIT and trained for 500 epochs.
@@
For TD-FT, we followed the approach in Section 4.2 by finetuning representations according to Chen et al., 2020b . Specifically, we first pretrained an unsupervised encoder following their pretraining setup (detailed below) then finetuned the encoder for 500 epochs from the first layer of the projection head. We used a batch size of 512, learning rate of 0.11, learning rate warmup and the LARS (You et al.,, 2017) optimiser with a momentum of 0.9.
For the US approach, we finetuned according to the pretraining setup in Chen et al., 2020b for CIFAR-10 as it was more reasonable for our datasets. Specifically, we pretrained for 500 epochs using a batch size of 512, learning rate of 2.26 which was linearly increased for the first 5% of epochs and subsequently decayed with the cosine decay schedule. We again used the LARS optimiser with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. We used a 1–layer projection head and a temperature of 0.2 for the contrastive loss. For our pretraining augmentations, we used random resized crop (area sampled uniformly from 8% to 100%) and color distortion (jitter and random grayscale) with strength 0.5.
@@
For all datasets, we performed data augmentations on our labeled set during the training of the semi–supervised encoders. We used random rotations between $20^{\circ}$ and $20^{\circ}$, random horizontal flips with probability 0.5, and random affine transformations with scale between 0.65 and 1.

### v2

@@ 4.2 A Representation Fine–Tuning Approach
The approach naturally starts with some initial representation defined by an encoder $g:X\to\mathbb{R}^{d}$. This can either be trained directly on the pool data using any powerful unsupervised representation learning technique—such as those based on contrastive learning (e.g. SimCLR (Chen et al., 2020a)), clustering (e.g. SwAV (Caron et al., 2020), DeepCluster (Caron et al., 2018)) or masked autoencoders (e.g. MAE (He et al., 2022))—or it can be taken as a fixed pretained encoder trained on some other data, such as ESM-3 for protein sequences (Hayes et al., 2025). In the latter case, the representation does not necessarily need to have been trained in an unsupervised manner itself, but it will inevitably not have information about the labels of the task at hand as these are yet to be collected. For our experiments we will use an encoder trained on the pool data using SimCLRv2 (Chen et al., 2020b).
Once initialised, we extend $g$ by adding a guidance classifier $c_{\omega}:\mathbb{R}^{d}\to[0,1]^{|\mathcal{Y}|}$ to its final layer, resulting in the model $g\circ c:X\to[0,1]^{|\mathcal{Y}|}$, which we train in a fully supervised manner using the acquired labels following Chen et al. (2020b). The final representations are taken from $g$ after this fine–tuning as before. For $c_{\omega}$, we follow Chen et al. (2020b) and use what is effectively a 1-layer neural network with a modest number of hidden units. For the same reasons discussed in §4.1, we do not use $c_{\omega}$ for our prediction head, again using a random forest in our experiments instead. Further algorithmic details on the precise approach used for the experiments and ablations with different prediction heads are given in the Appendix.
Table 1: Final test accuracy of different active learning methods on the F+MNIST, CIFAR-10+100 and CheXpert datasets. We report the mean $\pm 1$ standard error over 4 seeds. The full active learning curves are given in the Appendix.
@@
SIMILAR (Kothawade et al., 2021) $93.82\pm 0.18$ $30.87\pm 1.57$ $71.94\pm 0.47$
GALAXY (Zhang et al., 2022) $84.74\pm 0.74$ $55.28\pm 0.42$ $78.76\pm 0.35$
Cluster Margin (Citovsky et al., 2021) $94.24\pm 0.17$ $32.79\pm 0.45$ $75.41\pm 0.29$
US+EPIG (SimCLRv2, (Bickford Smith et al., 2024)) $98.53\pm 0.12$ $6.19\pm 0.42 77$$7.84\pm 0.28$
US+EPIG (VAE, (Bickford Smith et al., 2024)) $94.50\pm 0.34$ $30.47\pm 1.17$ $66.69\pm 0.56$
@@ 5 Related Work
Previous works on active learning with messy, uncurated pools have primarily focused on developing new acquisition strategies (Zhang et al., 2022; Kothawade et al., 2021; Zhang et al., 2023; Nuggehalli et al., 2023; Russakovsky et al., 2015), neglecting the importance of using a model that incorporates information from the unlabelled data. Indeed, they have mainly used fully supervised models, with some approaches initialising their model with weights pre–trained on ImageNet in a fully supervised fashion or from foundations models such as CLIP (Radford et al., 2021; Xie et al., 2024; Nuggehalli et al., 2023). We compare with these alternatives in §6.3 and find they significantly underperform compared to using representations trained on the pool.
@@
On the other hand, various work have considered using semi–supervised models in active learning (Burkhardt et al., 2018; Seo et al., 2022a; Osband et al., 2022a; Mittal et al., 2023; Sener & Savarese, 2017; Yehuda et al., 2022; Lüth et al., 2023), with several works showing supervised models significantly lagging behind semi–supervised ones in downstream performance (Bickford Smith et al., 2024; Hacohen et al., 2022; Yehuda et al., 2022). In general, there have been questions raised about the benefits of active learning when semi–supervised models are used: many acquisition strategies designed for fully supervised models have been shown to no longer provide reliable gains over random acquisition in the semi-supervised setting (Sener & Savarese, 2017; Yehuda et al., 2022; Lüth et al., 2023; Chan et al., 2021; Bhatnagar et al., 2020; Ebrahimi et al., 2020; Gao et al., 2020). However, Bickford Smith et al. (2024) recently showed that EPIG does reliably outperform random acquisition, as well as various other acquisition strategies, in this setting, using their unsupervised representation learning approach.
6 Experiments
We refer to the approaches introduced in §4.1, 4.2 as TD-SPLIT and TD-FT respectively. To validate them, we first compare with various baselines that have been specifically designed for, or shown strong performance on, the messy pool scenarios we are interested in (c.f. §3). As part of this, we include different variants of the unsupervised representation (US) approach of Bickford Smith et al. (2024). All experiments were run on an NVIDIA H100 80GB GPU.
Datasets: To construct datasets with redundant classes and class imbalance, we combined existing benchmarks. Our F+MNIST dataset uses the digits “5” and “6” from MNIST (Deng, 2012) as the target classes for active learning, while the entire FashionMNIST (Xiao et al., 2017) dataset is included as redundant data. CIFAR-10+100 uses the first five classes of CIFAR-10 (Krizhevsky, 2009) as its target classes, with the full CIFAR-100 dataset (Krizhevsky, 2009) serving as the redundant data.
We further use the CheXpert (Irvin et al., 2019) dataset as an example of a real–world dataset with redundant information and existing class imbalance. CheXpert comprises of chest X–rays taken from a variety of patients from different angles. We consider the binary classification task of identifying pleural effusion, i.e. fluid in the corner of the lungs. The redundant features here are the many anatomical and acquisition–related variations (e.g., bones, implants, soft tissue, scan artefacts) that are irrelevant to the diagnosis and generally represent larger image variations than the target–relevant information (Joy et al., 2021).
Representation learning: For all datasets, we use VAE (Kingma et al., 2013) and SimCLRv2 (Chen et al., 2020b) encoders, pairing them with TD–SPLIT and TD–FT respectively. We also include unsupervised variants following Bickford Smith et al. (2024) (noting that they also consider the same two encoder strategies). For all approaches, we adopt the learning rate, batch size, and optimizer from the original papers and train for 500 epochs.
Models: For F+MNIST and CheXpert, we use the Burgess encoder (Burgess et al., 2017) (and decoder) for our VAE–based representations and otherwise use a ResNet18 encoder (He et al., 2016). For CIFAR-10+100, we replace the ResNet18 with a ResNet50 and the Burgess encoder with a ResNet–VAE (Kingma et al., 2016).
Active learning: We run active learning for a budget of 500, 6000, and 10,000 labels for F+MNIST, CheXpert and CIFAR–10+100 respectively. We use batch acquisitions for all datasets, with a batch size of 10 for F+MNIST and 100 for the others. We use the “power” batch acquisition strategy from (Kirsch et al., 2021) with $\beta=4$ for F+MNIST and $\beta=8$ otherwise. We make this choice as this strategy is both highly scalable and has been shown to give performance comparable to more sophisticated batch acquisition strategies. We re–train our semi–supervised encoders every 5 acquisition rounds and ablate with different re–training periods in the Appendix.
@@ 6.1 Comparison with Existing Approaches
To highlight the effectiveness of our approach, we first compare against baselines designed for (or shown good performance on) problems with messy pools: SIMILAR (Kothawade et al., 2021), Cluster Margin (Citovsky et al., 2021), and GALAXY (Zhang et al., 2022), and the approach in Bickford Smith et al. (2024) which we refer to as US+EPIG. We also compare with random acquisition using the US, TD-SPLIT, and TD-FT approaches. We implement the baselines as in the original papers.
@@
Our TD-SPLIT method also shows strong performance relative to the baselines, albeit generally underperforming TD-FT. This is because it is using a much more lightweight encoder: it still consistently and comprehensively outperforms the analogous US+EPIG approach with VAE–based representations, thus again emphasising the importance of using task–driven representations.
@@
Figure 3: Test accuracy on F+MNIST using the US approach and our task–driven approach. Top and bottom row respectively show the results using VAE–based encoders and SimCLRv2 encoders. Solid line shows mean and shading $\pm 1$ standard error across 4 seeds.
@@
Figure 4: Test accuracy on CheXpert using the US approach and our task–driven approach. Top and bottom row respectively show the results using VAE–based encoders and SimCLRv2 encoders. Solid line shows mean and shading $\pm 1$ standard error across 4 seeds.
6.2 Acquisition Counts for Different Approaches
Table 2: Number of target classes that have been acquired at the end of active learning for different approaches on the F+MNIST and CIFAR-10+100 datasets. We report the mean $\pm 1$ standard error over 4 seeds. We omit the citations here for brevity.
Method F+MNIST CIFAR-10+100
SIMILAR $486\pm 6$ $920\pm 237$
GALAXY $118\pm 13$ $1806\pm 32$
Cluster Margin $32\pm 1$ $1287\pm 53$
US+EPIG (SimCLRv2) $119\pm 8$ $936\pm 35$
US+EPIG (VAE) $118\pm 4$ $529\pm 17$
Random $31\pm 2$ $609\pm 13$
TD-SPLIT (Ours) $141\pm 14$ $1651\pm 25$
TD-FT (Ours) $06\pm 38 22$$215\pm 38$
Table 2 shows the number of acquisitions that have been made for the target classes at the end of active learning for different approaches. We first note that using our approach instead of unsupervised representations improves the count of the target classes by a large margin. This suggests that the gains displayed in Table 5 are not merely from using a better model, but also from making better acquisitions (see also Table 6). Secondly, we note that the best performing method (TD-FT) does not always have the largest amount of target classes acquired. This suggests that the precise data point acquired is important, not just its class (Yang et al., 2023).
@@ 6.3 Pool-Based Task-Driven Representations Outperform Transfer Learning Based AL
A common setup in AL papers that deal with messy pools is a transfer–learning approach where a fully supervised model initialised from a large pretrained/foundation model is fine–tuned to the task (Zhang et al., 2022, 2023; Xie et al., 2024; Nuggehalli et al., 2023; Bommasani et al., 2021; Gupte et al., 2024). Our TD-FT framework can leverage these models by initialising the encoder, $g$, with their pretrained weights instead of using an unsupervised representation of the pool. We test this on the CIFAR-10+100 dataset by comparing our standard TD-FT method against initializing the encoder using both publicly available pre-trained models—namely ResNet-50 and ResNet-101 pre–trained with supervision on ImageNet (Russakovsky et al., 2015)—and foundation models—namely two variants of the self-supervised DinoV2 model (Oquab et al., 2023), pre–trained on a large, curated dataset of 142 million unlabelled images.
@@ 6.4 Alternative Acquisition Strategies
As an ablation to demonstrate the broader benefits of our task–driven approach with other acquisition strategies, we apply our TD-FT and TD-SPLIT approaches to BALD (Houlsby et al., 2011) and Confidence Sampling (Settles, 2009), comparing to analogous unsupervised (US) setups.
@@ A.1.1 Datasets
We used the F+MNIST as an example of a dataset with redundant classes and class imbalance by combining existing benchmarks. Specifically, we used the digits “5” and “6” from MNIST (Deng, 2012) as the target classes for active learning while the entire FashionMNIST dataset (Xiao et al., 2017) was included as the redundant data. We used an imbalance ratio of 10 for our pool, where the minority classes were chosen to be our target classes.
@@
We used the CIFAR-10+100 as another example of a dataset with redundant classes and class imbalance by combining existing benchmarks. Specifically, we used the first 5 classes of CIFAR-10 (Krizhevsky, 2009) as our target classes for active learning while the entire CIFAR-100 dataset (Krizhevsky, 2009) was included as the redundant data. We again used an imbalance ratio of 10 in our pool, where the minority classes were chosen to be our target classes.
@@
We used the CheXpert (Irvin et al., 2019) dataset as an example of a real–world dataset with redundant information and existing class imbalance. CheXpert comprises of chest X–rays taken from a variety of patients from different angles. We considered the binary classification task of identifying pleural effusion, i.e. fluid in the corner of the lungs. We filtered out observations with “NA” as the response for the pleural effusion task. The imbalance ratio in our pool was 2.5.
@@ A.1.2 Representation learning
For all datasets, we used VAE (Kingma et al., 2013) and SimCLRv2 encoders, pairing them with TD-SPLIT, TD-FT respectively. We also used their unsupervised variants for the US approach (Bickford Smith et al., 2024), where we first encoded our data points into the latent space of the unsupervised encoder, then performed active learning on the latent space with a prediction head. Below, we describe the details of the representation learning methods used.
@@
For TD-SPLIT, we followed the CCVAE (Joy et al., 2021) approach described in Section 4.1 and optimised objective (1). We chose $\alpha$ so that the labelled loss roughly matched the scale of the unlabelled loss. We performed the optimisation using stochastic gradient ascent where updates with the labelled and unlabelled data were conducted in separate batches. As the labelled dataset contained far fewer data points, we evenly spaced the labelled batches between the unlabelled ones. We trained our model for 500 epochs and, following Joy et al. (2021), we used a batch size of $200$ for both the unlabelled and labelled data, the Adam optimiser (Kingma & Ba, 2014) and a learning rate of $2\times 10^{-4}$. Table 3 shows the sizes for $z$, $z_{c}$ and values of $\alpha$.
For the US approach, we followed Burgess et al. (2017) and optimised their ELBO objective with $\beta=1$ (this amounts to the standard VAE objective). We used the Adam optimiser, a learning rate of $5\times 10^{-4}$, a batch size of $200$ and KL annealing as in Burgess et al. (2017). For all datasets, we used the same latent dimensions as for TD-SPLIT and trained for 500 epochs.
@@
For TD-FT, we followed the approach in Section 4.2 by finetuning representations according to Chen et al. (2020b). Specifically, we first pretrained an unsupervised encoder following their pretraining setup (detailed below) then finetuned the encoder for 500 epochs from the first layer of the projection head. We used a batch size of 512, learning rate of 0.11, learning rate warmup and the LARS (You et al., 2017) optimiser with a momentum of 0.9.
For the US approach, we finetuned according to the pretraining setup in Chen et al. (2020b) for CIFAR-10 as it was more reasonable for our datasets. Specifically, we pretrained for 500 epochs using a batch size of 512, learning rate of 2.26 which was linearly increased for the first 5% of epochs and subsequently decayed with the cosine decay schedule. We again used the LARS optimiser with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. We used a 1–layer projection head and a temperature of 0.2 for the contrastive loss. For our pretraining augmentations, we used random resized crop (area sampled uniformly from 8% to 100%) and color distortion (jitter and random grayscale) with strength 0.5.
@@
For all datasets, we performed data augmentations on our labelled set during the training of the semi–supervised encoders. We used random rotations between $20^{\circ}$ and $20^{\circ}$, random horizontal flips with probability 0.5, and random affine transformations with scale between 0.65 and 1.